OpenStudy (anonymous):

the radius of a pizza is 20 cm, what is the central angle in degrees that gives one person 30 cm of crust?

OpenStudy (anonymous):

You should use this: \[s=\theta r\] How?

OpenStudy (anonymous):

i dont know how?

OpenStudy (anonymous):

r is the radius, theta the angle, s the length of the crust. How can we use that?

OpenStudy (anonymous):

i did thatt way but i couldnt get the answer

OpenStudy (nali):

use this formula instead minor arc = (central angle / 360) (circumference) the minor arc is the 30cm = (central angle/ 360) (circumference of the pizza)

OpenStudy (nali):

do u get it so far?

OpenStudy (anonymous):

no i didnt

OpenStudy (nali):

do u still need help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have four answers but i dont know which is corect/

OpenStudy (anonymous):

may i know your answer

OpenStudy (nali):

ok do u at least get the equation

OpenStudy (anonymous):

no its confucing

OpenStudy (anonymous):

I got this: \[\theta = \frac{30}{20}\] \[\theta=\frac{3}{2}\] theta in radians

OpenStudy (anonymous):

yes i got taht

OpenStudy (anonymous):

And that's wrong?

OpenStudy (anonymous):

Maybe we have to write it in degrees: 2/3 rad=38.2 °

OpenStudy (anonymous):

a.120 b. 3 c. 4 d. 86 these are the four answers that i have to choose from

OpenStudy (anonymous):

120 b. 35 c. 43 d. 86

OpenStudy (nali):

i know the answer if u want it

OpenStudy (anonymous):

yes can i have please

OpenStudy (nali):

it is 86 so d

OpenStudy (anonymous):

are you sure. because this is my last online class test.

OpenStudy (nali):

yes i am sure

OpenStudy (anonymous):

thank you :)

OpenStudy (nali):

ur welcome sorry to make u wait all this long.. :))